# Ergodicity for locally monotone stochastic evolution equations with Lévy noise

### v1

@@ 2. The model and main results
We are interested in SPDEs with locally monotone drift and additive Lévy noise, which consists of a Gaussian part and an independent Poisson part with small jumps. Let $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space satisfying the standard conditions. Let $H$ be a separable Hilbert space. Let $V$ be a reflexive Banach space embedded linearly, densely, and continuously to $H$. We would like to point out that we do not assume compactness of the embedding, even if the embedding may be compact in several examples. We consider the unique strong solution $(X_{t})_{t\geq 0}$ of
(2.1) $\mathrm{d}X_{t}=A(X_{t})\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}=x\in H,$
where $(W_{t})_{t\geq 0}$ is a $H$-valued cylindrical Wiener process on $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t})_{t\geq 0},\mathbb{P})$, see [21]. The locally monotone nonlinear drift operator $A:V\to V^{\ast}$ and the bounded linear operator $B:H\to H$ satisfy the hypotheses in Section 3, in particular, it has finite Hilbert-Schmidt norm $\|B\|_{2}:=\|B\|_{L_{2}(H)}$. Assume that $(Z,{\mathcal{Z}},\pi)$ is a $\sigma$-finite measure space, where $\pi$ is the intensity measure of a stationary Poisson point process on $Z$. Denote by $D^{c}$ (representing small jumps) the complement of a fixed measurable set $D\in{\mathcal{Z}}$ (representing large jumps), where we assume that
(2.2) $\pi(D)<\infty.$
Denote the compensated Poisson random measure $\widetilde{N}:(0,\infty)\times\Omega\times{\mathcal{Z}}\to\mathbb{R}\cup\{-\infty,\infty\}$ by
(2.3) $\widetilde{N}((0,t]\times U)=N((0,t]\times U)-t\,\pi(U),\quad t>0,\quad U\in{\mathcal{Z}},$
where $N:[0,T]\times\Omega\times{\mathcal{Z}}\to[0,\infty]$ is a Poisson random measure with intensity $\pi$, in particular,
(2.4) $\mathbb{E}[N((0,t]\times U)]=t\,\pi(U),\quad t>0,\quad U\in{\mathcal{Z}},$
see [74, 14, 51]. Let $G:Z\to H$ be a strongly ${\mathcal{Z}}$-measurable function satisfying Hypothesis (F). For the rigorous definition of the stochastic integrals, we refer to [14, 21, 71, 11].
@@ 2.1. Main results
Our main result is the following abstract ergodicity result for SPDEs of the type (2.1) with locally monotone drift with degenerate and spatially regular Lévy noise. See Subsection 3.4 for the precise definitions of the terminology.
@@
Assume that Hypotheses (A)(B)(C)(D)(E)(F) and (G) given in Section 3 hold true. Then, the stochastically continuous Markovian Feller semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) satisfies the $e$-property in $H$ and is weak-mean ergodic. Moreover, it admits a unique invariant probability measure $\mu_{\ast}$ on $(H,{\mathcal{B}}(H))$ that admits finite $(\alpha+\beta)$-moments in $H$, where $\alpha\geq 2$ and $\beta\geq 0$ are as in Hypothesis (C). If $\beta=0$, $\mu_{\ast}$ admits finite $\alpha$-moments in $V$.
Under the additional assumption that $V\subset H$ is a compact embedding (which we do not assume), Hypothesis (B) already guarantees that the existence of at least one invariant measure, which can be proved by the classical method of Krylov and Bogoliubov, see [20, Theorem 3.1.1].
In Section 4, we discuss the following examples. First, for the sake of completeness, we consider the stochastic heat equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, with Dirichlet boundary conditions on a Lipschitz boundary,
(2.5) $\mathrm{d}X_{t}=\Delta X_{t}\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}\in L^{2}({\mathcal{O}}).$
We report the following result, which is well-known for the case of Gaussian noise, see [8, 20]. See [68] for the the case of degenerate Lévy noise.
@@
Assume that Hypotheses (F) and (G) hold true. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic heat equation (2.5) with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite second moments in the stronger space $V$ in all spatial dimensions.
Furthermore, we consider the semilinear stochastic equation on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\{1,2,3\}$, $\nu>0$, with Lipschitz boundary,
(2.6) $\mathrm{d}X_{t}=\left(\nu\Delta X_{t}+\mathbf{f}(X_{t})\cdot\nabla X_{t}+g(X_{t})\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad X_{0}\in L^{2}({\mathcal{O}}).$
For $\mathbf{f}:=(f_{1},\ldots,f_{d})^{\textup{t}}$, where $v^{\textup{t}}$ denotes the transpose of $v$, we assume that $f_{i}\in\operatorname{Lip}_{b}(\mathbb{R})$, $i=1,\ldots,d$ for $d=1,2$. If $d=3$, we consider the case $f_{i}:=b_{i}\in L^{\infty}({\mathcal{O}})$, $i=1,\ldots,3$. Let $c_{0}>0$ be the embedding constant of the embedding $V\subset H$, see Fact 3.1 below. We assume that $g:\mathbb{R}\to\mathbb{R}$ is continuous with $g(0)=0$, and that there exist non-negative constants $s,C,K,k$ and $c\in(0,\nu c_{0}^{2})$, such that
@@
such that for $d=1,2$, we assume that $s\leq 2$, and for $d=3$, we assume that $s\leq\frac{4}{3}$. For $d=1$, assume that $c<\nu c_{0}^{2}$. For $d=2,3$, assume that
@@
Assume that $\mathbf{f}$ and $g$ satisfy (2.7), (2.8), (2.9), (2.10) and (2.11) and assume that Hypotheses (F) and (G) for the noise are valid. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic semilinear equations of the form (2.6) for $d\leq 3$, with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite second $V$-moments for $d=1,2$, and finite $\frac{16}{3}$th moments for $d=3$. In particular, this holds for the stochastic Burgers equation in 1D for $f_{1}(x)=x$.
For the stochastic incompressible 2D Navier-Stokes equations on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{2}$, $d\in\mathbb{N}$, with Dirichlet boundary conditions on a Lipschitz boundary $\partial{\mathcal{O}}$, with viscosity $\nu>0$,
(2.12) $\mathrm{d}X_{t}=\left(\nu\mathbf{P}\Delta X_{t}-\mathbf{P}\left[(X_{t}\cdot\nabla)X_{t}\right]\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad\mathbf{P}(X_{0})\in L^{2}({\mathcal{O}};\mathbb{R}^{2}),$
where $\mathbf{P}$ denotes the Helmholtz-Leray projection on the solenoidal fields, we get the following result, which is, to our knowledge, novel for degenerate noise.
@@
Assume that Hypotheses (F) and (G) for the noise hold true. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic incompressible 2D Navier-Stokes equations (2.12) with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite fourth moments.
As a generalization, one can consider the velocity field of a viscous and incompressible non-Newtonian fluid perturbed by Lévy noise with Dirichlet boundary conditions on a bounded domain ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Lipschitz boundary $\partial{\mathcal{O}}$, $d\geq 2$. Let $p>2$, $\nu>0$, and $u:{\mathcal{O}}\to\mathbb{R}^{d}$. Define
@@
(2.13) $\mathrm{d}X_{t}=\left(\nu\mathbf{P}(\operatorname{div}(\tau(X_{t})))-\mathbf{P}\left[(X_{t}\cdot\nabla)X_{t}\right]\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t}+\int_{D^{c}}G(z)\,\widetilde{N}(\mathrm{d}t,\mathrm{d}z),\quad\mathbf{P}(X_{0})\in L^{2}({\mathcal{O}};\mathbb{R}^{d}).$
@@
Assume $p\geq\frac{2+d}{2}\vee 2$, and that Hypotheses (F) and (G) hold true. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic shear thickening power-law fluid equations (2.13) with degenerate additive Lévy noise possesses a unique invariant measure $\mu_{\ast}$ with finite $\frac{2p}{2p-d}$th moments.
@@ 3. Preliminaries and hypotheses
The following hypotheses are modifications and partial extensions, needed for our ergodicity result, of the hypotheses for the variational well-posedness result for SPDEs with locally monotone drift in finite time from [55, 56] for Gaussian noise, which has been extended to Lévy noise in [14]. See also [36, 65, 3, 50, 69, 1, 2] for further extensions, which cover examples that are out of the scope of this paper; for instance the 3D stochastic tamed Navier-Stokes equation, the stochastic $p$-Laplace equation, the stochastic Allen-Cahn equation, and the stochastic Cahn-Hilliard equation, where the drift of the latter SPDE satisfies bounds which have been called fully locally monotone in [69].
@@ 3.1. Hypotheses on the drift operator
$\mathbb{R}\ni\lambda\mapsto\langle A(x+\lambda y),z\rangle\quad\textrm{ is continuous.}$.
@@
Hypothesis (C) (Local monotonicity).
@@
$2\langle A(x)-A(y),x-y\rangle\leq[-\delta_{2}+\rho(y)]\|x-y\|^{2}_{H}\quad\textrm{ for all }\quad x,y\in V,$
where $\alpha\geq 2$ is given in Hypothesis (B) and $\rho:V\to[0,\infty)$ is a measurable function, locally bounded in $V$, satisfying
$0\leq\rho(y)\leq C_{2}\|y\|^{\alpha}_{V}\|y\|^{\beta}_{H}\quad\textrm{ for any}\quad y\in V.$
@@
Note that above, especially Hypothesis (C) and Hypothesis (E) differ from the standard assumptions for locally monotone drift SPDEs [56]. To obtain global coupling at infinity (the so-called $e$-property, see Definition 3.7 below), we need to obtain a priori estimates independent of the terminal time $T>0$. To get this, we need to assume $\delta_{2}>0$, and we cannot permit constant positive perturbations of the drift. Essentially, this means that our drift has a strongly dissipative part. Hypothesis (E) is a cone condition that is of technical nature, but is easy to verify for many examples where the growth of the lower order nonlinearities can be controlled by the potential of the leading order term, see the examples in Section 4 below. Hypothesis (E) can for instance be verified with the help of the conditions of Lemma 4.1 and Lemma 4.2 below. Note that we need Hypothesis (E) only in the proof of Lemma 5.8 below.

### v4

@@ 2. The model and main results
We are interested in SPDEs with locally monotone drift and additive Wiener noise. Let $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t})_{t\geq 0},\mathbb{P})$ be a filtered probability space satisfying the standard conditions. Let $H$ be a separable Hilbert space. Let $V$ be a reflexive Banach space embedded linearly, densely, and continuously to $H$. We would like to point out that we generally do not assume compactness of the embedding which enables us to consider certain unbounded spatial domains, namely, those of finite width.
We consider the unique strong solution $(X_{t})_{t\geq 0}$ of
(2.1) $\mathrm{d}X_{t}=A(X_{t})\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}=x\in H,$
where $(W_{t})_{t\geq 0}$ is a $U$-valued cylindrical Wiener process on $(\Omega,{\mathcal{F}},({\mathcal{F}}_{t})_{t\geq 0},\mathbb{P})$ for some separable Hilbert space $U$, see [25]. The locally monotone nonlinear drift operator $A:V\to V^{\ast}$ and the bounded linear operator $B:U\to H$ satisfy the hypotheses in Section 3, in particular, it has finite Hilbert-Schmidt norm $\|B\|_{L_{2}(U,H)}<\infty$. Note that $B$ can be equal to zero.
@@ 2.1. Main results
Our main result is the following abstract ergodicity result for SPDEs of the type (2.1) with locally monotone drift with degenerate and spatially regular Wiener noise. See Subsection 3.5 for the precise definitions of the terminology.
@@
Assume that Hypotheses (A)(B)(C)(D)(E)(F)(G), and (H) given in Section 3 hold true. Then, the stochastically continuous Markovian Feller semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) satisfies the $e$-property in $H$ and is weak-mean ergodic. Moreover, it admits a unique invariant probability measure $\mu_{\ast}$ on $(H,{\mathcal{B}}(H))$ that admits finite $(\alpha+\beta)$-moments in $H$, where $\alpha\geq 2$ and $\beta\geq 0$ are as in Hypothesis (C). If $\beta=0$, $\mu_{\ast}$ admits finite $\alpha$-moments in $V$.
Proof.
See Section 5. ∎
Under the additional assumption that $V\subset H$ is a compact embedding (which we do not assume), Hypothesis (B) already guarantees that the existence of at least one invariant measure, which can be proved by the classical method of Krylov and Bogoliubov, see [24, Theorem 3.1.1].
See Section 6 for the terminology of $\varepsilon$-mixing times.
@@
Assume that Hypotheses (A)(B)(C)(D)(E)(F)(G)(H), and (I) given in Section 3 hold true. Then, the stochastically continuous Markovian Feller semigroup $(P_{t})_{t\geq 0}$ associated to (2.1) admits an exponentially ergodic unique invariant probability measure $\mu_{\ast}$ on $(H,{\mathcal{B}}(H))$ such that the $\varepsilon$-mixing time of $\mu_{\ast}$ in Wasserstein-2-distance is bounded above. See Theorem 6.2 for the explicit bound.
Proof.
See Section 6.1. ∎
In Section 4, we discuss the following examples. First, we consider the stochastic heat equation on a domain of finite width [^1] A domain ${\mathcal{O}}\subset\mathbb{R}^{d}$ of finite width fits by definition between two parallel (d1)(d-1)-dimensional hyperplanes, see [1]. Sometimes it is referred to as a domain which is bounded in one direction. ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d\in\mathbb{N}$, $\nu>0$, with Dirichlet boundary conditions on a Lipschitz boundary,
(2.2) $\mathrm{d}X_{t}=\nu\Delta X_{t}\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}\in L^{2}({\mathcal{O}}).$
We report the following result, compare also with [7, 24].
Theorem 2.3.
Assume that $B\in L_{2}(U,W_{0}^{1,2}({\mathcal{O}}))$. Denote by $c_{0}$ the the inverse Poincaré constant of $\mathcal{O}$}, and assume that for some $\lambda\in(0,1)$,
$\|B\|^{2}_{L_{2}(U,H)}\leq\lambda\nu c_{0}^{2}.$
Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic heat equation (2.2) with degenerate additive Wiener noise is exponentially ergodic and exponentially mixing and possesses a unique invariant measure $\mu_{\ast}$ with finite second moments in the stronger space $V$ in all spatial dimensions. We have the following upper bound for the $\varepsilon$-mixing time
$\tau^{x}_{\textup{{mix}}}(\varepsilon)\leq\frac{1}{\nu c_{0}^{2}}\left[\log\left(\|x\|_{H}+\frac{\|B\|_{L_{2}(U,H)}}{\sqrt{2(1-\lambda)\nu}c_{0}}\right)+\log\left(\frac{1}{\varepsilon}\right)\right].$
Proof.
See Subsection 4.1. ∎
Furthermore, we consider the semilinear stochastic equation on a domain of finite width ${\mathcal{O}}\subset\mathbb{R}^{d}$, $d=1,2$, $\nu>0$, with Lipschitz boundary,
(2.3) $\mathrm{d}X_{t}=\left(\nu\Delta X_{t}+\mathbf{f}(X_{t})\cdot\nabla X_{t}+g(X_{t})\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}\in L^{2}({\mathcal{O}}).$
For $\mathbf{f}:=(f_{1},\ldots,f_{d})^{\textup{t}}$, where $v^{\textup{t}}$ denotes the transpose of $v$, we assume that $f(x)=x$ in $d=1$ or that $f_{i}\in\operatorname{Lip}_{b}(\mathbb{R})$, $i=1,\ldots,d$ for $d=1,2$. Let $c_{0}>0$ be the inverse Poincaré constant of $\mathcal{O}$}. We assume that $g:\mathbb{R}\to\mathbb{R}$ is continuous with $g(0)=0$, and that there exist non-negative constants $C,c,s,K,k$, such that
@@
such that $s\leq 2$. For $d=1$, assume that $c<\nu c_{0}^{2}$. For $d=2$, assume that
@@
Assume that $\mathbf{f}$ and $g$ satisfy (2.4), (2.5), (2.6), (2.7) and (2.8) and assume that $B\in L_{2}(U,W_{0}^{1,2}({\mathcal{O}}))$. Assume that Hypotheses (H) and (I) hold. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic semilinear equations of the form (2.3) for $d=1,2$, with degenerate additive Wiener noise is exponentially mixing and exponentially ergodic and possesses a unique invariant measure $\mu_{\ast}$ with finite second $V$-moments. In particular, this holds for the stochastic Burgers equation in 1D for $f_{1}(x)=x$.
Proof.
See Subsection 4.2. ∎
For the stochastic incompressible 2D Navier-Stokes equations on a domain of finite width ${\mathcal{O}}\subset\mathbb{R}^{2}$ with no-slip (that is, Dirichlet) boundary conditions on a Lipschitz boundary $\partial{\mathcal{O}}$, with viscosity $\nu>0$,
(2.9) $\mathrm{d}X_{t}=\left(\nu\mathbf{P}\Delta X_{t}-\mathbf{P}\left[(X_{t}\cdot\nabla)X_{t}\right]\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}\in L^{2}_{\text{sol}}({\mathcal{O}};\mathbb{R}^{2}),$
where $\mathbf{P}$ denotes the Helmholtz-Leray projection on the solenoidal fields, and $L^{2}_{\operatorname{sol}}({\mathcal{O}}):=\mathbf{P}(L^{2}({\mathcal{O}}))$, $W^{1,2}_{0,\,\operatorname{sol}}({\mathcal{O}}):=\mathbf{P}(W^{1,2}_{0}({\mathcal{O}}))$. Let $c_{0}>0$ be the inverse Poincaré constant of $\mathcal{O}$}. We get the following result, which is, to our knowledge, novel for degenerate noise.
@@
Assume that $B\in L_{2}(U,W^{1,2}_{0,\,\operatorname{sol}}({\mathcal{O}}))$. Assume that for some $\lambda\in(0,1)$,
$\|B\|^{2}_{L_{2}(U,H)}\leq\frac{1}{4}\lambda\nu^{3}c_{0}^{2},$
and
$\|B\|^{2}_{L_{2}(U,H)}\leq(1-\lambda)\nu c_{0}^{2}.$
Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic incompressible 2D Navier-Stokes equations (2.9) with degenerate additive Wiener noise possesses a unique invariant measure $\mu_{\ast}$ with second moments in $V$.
Furthermore, if there exists $\gamma\in(0,\nu c_{0}^{2}]$ and $\lambda\in(0,1)$ such that
$\|B\|^{2}_{L_{2}(U,H)}\leq\frac{1}{4}\lambda(\nu^{3}c_{0}^{2}-\nu^{2}\gamma),$
and
$\|B\|^{2}_{L_{2}(U,H)}\leq(1-\lambda)\nu c_{0}^{2},$
then $(P_{t})_{t\geq 0}$ is exponentially mixing and exponentially ergodic, and we have the following upper bound for the $\varepsilon$-mixing time
$\tau^{x}_{\textup{{mix}}}(\varepsilon)\leq\frac{2}{\gamma}\left[\frac{1}{\lambda\nu^{2}}\|x\|^{2}_{H}+\log\left(\|x\|_{H}+\frac{\|B\|^{2}_{L_{2}(U,H)}}{\sqrt{2\lambda\nu}c_{0}}\right)+\log\left(\frac{1}{\varepsilon}\right)\right].$
Proof.
See Subsection 4.3. ∎
As a generalization, one can consider the velocity field of a viscous and incompressible non-Newtonian fluid perturbed by Wiener noise with Dirichlet boundary conditions on a domain of finite width ${\mathcal{O}}\subset\mathbb{R}^{d}$ with Lipschitz boundary $\partial{\mathcal{O}}$, $d\geq 2$. Let $p=1+\frac{d}{2}$, $\nu>0$, and $u:{\mathcal{O}}\to\mathbb{R}^{d}$. Define
@@
(2.10) $\mathrm{d}X_{t}=\left(\mathbf{P}(\operatorname{div}(\tau(X_{t})))-\mathbf{P}\left[(X_{t}\cdot\nabla)X_{t}\right]\right)\,\mathrm{d}t+B\,\mathrm{d}W_{t},\quad X_{0}\in L^{2}_{\text{sol}}({\mathcal{O}};\mathbb{R}^{2}).$
@@
Assume $p=1+\frac{d}{2}$ for $d\geq 2$, and that Hypotheses (F)(G)(H) hold true. Then the Markovian Feller semigroup $(P_{t})_{t\geq 0}$ of the stochastic shear thickening power-law fluid equations (2.10) with degenerate additive Wiener noise possesses a unique invariant measure $\mu_{\ast}$ with finite $(1+\frac{d}{2})$th moments. If, additionally, Hypothesis (I) holds, then $(P_{t})_{t\geq 0}$ is exponentially ergodic and exponentially mixing, with a quantitative upper bound for the $\varepsilon$-mixing time given in Theorem 6.2.
Proof.
See Subsection 4.4. ∎
It is easy to see, that the situation $p=d=2$, that is, the 2D Navier-Stokes case, as well as e.g. $p=\frac{5}{2}$ and $d=3$ are covered in the assumptions of Theorem 2.6. However, the 3D Navier-Stokes case (that is, $p=2$ and $d=3$) is not covered. Compare also with the results of [16] and [85].
@@ 3. Preliminaries and hypotheses
The following hypotheses are modifications and partial extensions, needed for our ergodicity result, of the hypotheses for the variational well-posedness result for SPDEs with locally monotone drift in finite time from [69, 70] for Gaussian noise, which has been extended to Lévy noise in [16]. See also [46, 80, 4, 63, 83, 2, 3] for further extensions, which cover examples that are out of the scope of this paper; for instance the 3D stochastic tamed Navier-Stokes equations, the stochastic $p$-Laplace equation, the stochastic Allen-Cahn equation, and the stochastic Cahn-Hilliard equation, where the drift of the latter SPDE satisfies bounds which have been called fully locally monotone in [83].
@@ 3.1. Hypotheses on the drift operator
$\mathbb{R}\ni\lambda\mapsto\langle A(x+\lambda y),z\rangle\quad\textrm{ is continuous.}$
@@
Hypothesis (C) (Full local monotonicity).
@@
$2\langle A(x)-A(y),x-y\rangle\leq[-\delta_{2}+\eta(x)+\rho(y)]\|x-y\|^{2}_{H}\quad\textrm{ for all }\quad x,y\in V,$
where $\alpha\geq 2$ is given in Hypothesis (B), and $\rho,\eta:V\to[0,\infty)$ are locally bounded, measurable functions, satisfying
$0\leq\eta(x)+\rho(x)\leq C_{2}\|x\|^{\alpha}_{V}\|x\|^{\beta}_{H}\quad\textrm{ for any }\quad x\in V.$
If both $\eta$ and $\rho$ are non-trivial, we also assume that the embedding $V\subset H$ is compact.
If both $\eta$ and $\rho$ in Hypothesis (C) are non-trivial, we say that we are in the fully local monotone case, see [83]. If Hypothesis (C) holds with either $\eta\equiv 0$ or $\rho\equiv 0$, we say that we are in the locally monotone case, see [70]. If Hypothesis (C) holds with $\eta=\rho\equiv 0$, we say that we are in the strongly monotone case, see [58].
@@
Note that above, especially Hypothesis (C) and Hypothesis (E) differ from the standard assumptions for locally monotone drift SPDEs [70]. To obtain global coupling at infinity (the $e$-property, see Definition 3.9 below), we need to obtain a priori estimates independent of the terminal time $T>0$. To get this, we need to assume $\delta_{2}>0$, and we cannot permit constant positive perturbations of the drift. Essentially, this means that our drift has a strongly dissipative part. Hypothesis (E) is a cone condition that is of technical nature, but is easy to verify for many examples where the growth of the lower order nonlinearities can be controlled by the potential of the leading order term, see the examples in Section 4 below. Hypothesis (E) can for instance be verified with the help of the conditions of Lemma 4.1 and Lemma 4.2 below. Note that we need Hypothesis (E) only in the proof of Lemma 5.6 below.